OpenStudy (anonymous):

It's been a while since I did 'no solution/infinite' solution problems, could somebody explain to me the basics please? Thank you! Here's one of the problems on my packet as an example: 3x-1+4(x) a. x=1 b. infinite solutions c. no solution c.x=-1

OpenStudy (anonymous):

the question is not an equation. post it in the form of an equation.

OpenStudy (anonymous):

whoops, the complete problem is 3x-1+4(x)=-2 sorry :/

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

so, does that mean no solution? why is there no solution?

OpenStudy (anonymous):

no, it doesn't mean that, otherwise I would have put "no solution" as the answer. the problem does have a solution, and it is none of the above.

OpenStudy (anonymous):

are you sure? my teacher listed the above answers as, well, the /only/ answers D:

OpenStudy (anonymous):

ooh, i see. should be -3x D: wow, i'm sorry. i'm really distracted.

OpenStudy (anonymous):

2x+4y=7 4x+8y=14 If you try to find points of intersection you will realize that they are same line thus there are infinte points of intersection 2x+4y=7 4x+8y=11 You will find that above lines are parellel thus there are no soloution

OpenStudy (anonymous):

-3x-1+4x = -2 4x -3x = -2+1 x = -1

OpenStudy (anonymous):

how do i know if there are infinite solutions, though? They only gave me one equation; there's no second one to line it up with.

OpenStudy (anonymous):

a "solution" is a point of intersection between two curves.

OpenStudy (anonymous):

what are the two curves in the problem? sorry for all the questions, history's always been my strong point , not algebra :/

OpenStudy (anonymous):

this is not a curve